# Bosch China Investment Ltd

### v4

@@ Abstract
Message hiding, a technique that conceals secret message bits within a cover image, aims to achieve an optimal balance among message capacity, recovery accuracy, and imperceptibility. While Convolutional Neural Networks (CNNs) have notably improved message capacity and imperceptibility, obtaining high recovery accuracy remains challenging. This arises because, in contrast to image data, the sequential arrangement of message bits is imperative, but convolutional operations encounter inherent difficulties in preserving this critical order. In addition, CNNs also struggle to effectively address the large discrepancy between these two uncorrelated modalities. To address this, we propose StegaFormer, an innovative MLP-based framework designed to preserve bit order and enable global fusion between modalities. Specifically, StegaFormer incorporates three crucial components: Order-Preserving Message Encoder (OPME), Order-Preserving Message Decoder (OPMD), and Global Message-Image Fusion (GMIF). OPME and OPMD aim to preserve the order of message bits by segmenting the entire sequence into equal-length segments and incorporating sequential information during encoding and decoding. Meanwhile, GMIF employs a cross-modality fusion mechanism to effectively fuse the features from the two uncorrelated modalities. Experimental results on the COCO and DIV2K datasets demonstrate that StegaFormer significantly outperforms existing state-of-the-art methods in terms of recovery accuracy, message capacity, and imperceptibility. Our code is released in https://github.com/boschresearch/Stegaformer.
@@ 1 Introduction
Image steganography is a technique that conceals confidential information within a publicly accessible image [Mielikainen(2006), Pevnỳ et al.(2010)Pevnỳ, Filler, and Bas, Zhang et al.(2019a)Zhang, Cuesta-Infante, and Veeramachaneni, Tan et al.(2021)Tan, Liao, Liu, Cao, and Jiang].In this study, we focus on message hiding. It has a wide range of applications such as copy-right protection, light field messaging [Wengrowski and Dana(2019)], virtual reality and augmented reality applications [Tancik et al.(2020)Tancik, Mildenhall, and Ng] and face anonymization [Kishore et al.(2021)Kishore, Chen, Wang, Li, and Weinberger]. The goal of it is embedding secret message bits within a cover image to produce a new image (called stego image). Ensuring a balanced trade-off among message capacity, recovery accuracy, and imperceptibility is the core task. More concretely, message capacity is measured by bits per pixel (BPP). Recovery accuracy indicates the message fidelity after being processed by these methods. Imperceptibility is minimizing the perceptible differences between the cover image and the stego image. It’s usually measured by the image quality metrics such as peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM) [Wang et al.(2004)Wang, Bovik, Sheikh, and Simoncelli].
@@
Figure 1: The illustration of our proposed methods. (a) CNN-based message encoding that disorders the message bits in message encoding and (b) our proposed message encoding: Order-Preserving Message Encoder (OPME) that preserves the order of message bits by segmenting and MLP-based encoding. (c) CNN-based message-image feature fusion and (d) our proposed message-image fusion using Global Message-Image Fusion (GMIF).
@@
Traditional message hiding methods often rely on heuristic approaches, such as the Least Significant Bit method [Mielikainen(2006)], which modifies individual pixels of a digital image to conceal message bits. Recently, with the advancements of deep learning, Convolutional Neural Network (CNN) based methods have significantly improved the message capacity and imperceptibility [Zhang et al.(2019a)Zhang, Cuesta-Infante, and Veeramachaneni, Yu(2020), Tan et al.(2021)Tan, Liao, Liu, Cao, and Jiang, Wei et al.(2022)Wei, Li, Zhang, Luo, Qian, and Zhou, Yin et al.(2023)Yin, Wu, Wang, Lu, Zhou, and Huang]. These methods involve aligning and concatenating message bits with the cover image in a spatial manner while utilizing CNNs to seamlessly integrate the message and image.
@@
However, CNN-based approaches often yield suboptimal results in message encoding and recovery, particularly when employing convolutional operations to encode multiple layers of message bits (message capacity > 1 BPP) as illustrated in Fig. 1(a). We have identified two main limitations in current CNN-based methods. Firstly, the indiscriminate use of convolutional kernels to encode message bits disrupts the sequential order of the original message bits. Secondly, the element-wise concatenation of message and image features, followed by convolutional operations, underestimates the substantial disparity between these two uncorrelated modalities, resulting in inadequate fusion as shown in Fig. 1(c).
@@
To address these challenges, as shown in Fig. 1(b), we leverage on MLP to design a more appropriate method for message encoding. The MLP-based message encoding phase can maintain the local order of small message segments. And the Multi-Head Self-Attention (MHSA) [Vaswani et al.(2017b)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez, Kaiser, and Polosukhin] and positional embeddings (PEs) help to encode the message features while explicitly adding global order information to the encoded message features. In this context, “local” refers to the order within a message segment, while “global” refers to the sequential order of the message segments. For message and image fusion, as shown in Fig. 1(d), we use Windowed Multi-Head Self-Attention Layer (W-MSA) to address the limitation of receptive fields in the CNN-based structure, enabling global interactions between the uncorrelated message and image features.
Our proposed approach, named StegaFormer, consists of three modules for enhanced message encoding, decoding, and message-image fusion, respectively. We design the first two modules, Order-Preserving Message Encoder (OPME) and Order-Preserving Message Decoder (OPMD), to effectively handle message segments while preserving their order. The OPMD adopts a symmetric structure and methodology with OPME for decoding. Additionally, OPME and OPMD fully leverage sequential order by incorporating PEs into the message features and globally encoding and decoding the features using MHSA. The last module, Global Message-Image Fusion(GMIF), employs Windowed Multi-Head Self-Attention Layer (W-MSA) [Liu et al.(2022)Liu, Hu, Lin, Yao, Xie, Wei, Ning, Cao, Zhang, Dong, et al.] to globally fuse the information from the secret message and cover image into the stego image, greatly minimizing perceptible artifacts in the stego image.
Our contributions are summarized as follows: (1) We are the first to explore the limitations of message hiding using CNN-based architecture; consequently, we transfer to an MLP-based approach design. (2) We introduce two novel modules OPME & OPMD, which enhance the message encoding and decoding by incorporating sequential order into the message. (3) We design GMIF, which leverages the global interaction of W-MSA to effectively fuse secret message and cover image features to stego image features. (4) Experimental results on real-world datasets demonstrate the superiority of our framework over existing state-of-the-art methods in terms of accuracy, capacity, and imperceptibility.
2 Related Work
@@
Recently, CNN-based methods for message hiding have significantly improved capacity and gained considerable attention [Zhang et al.(2019a)Zhang, Cuesta-Infante, and Veeramachaneni, Tan et al.(2021)Tan, Liao, Liu, Cao, and Jiang, Chen et al.(2022)Chen, Kishore, and Weinberger, Yin et al.(2023)Yin, Wu, Wang, Lu, Zhou, and Huang]. For instance, SteganoGAN [Zhang et al.(2019a)Zhang, Cuesta-Infante, and Veeramachaneni] introduced 3D message tensor alignment along spatial dimensions with the image tensor, and ChatGAN [Tan et al.(2021)Tan, Liao, Liu, Cao, and Jiang] incorporated channel attention to enhance accuracy and imperceptibility while maintaining high capacity. SFTN [Yin et al.(2023)Yin, Wu, Wang, Lu, Zhou, and Huang] was proposed to address rounding errors before generating the stego image. However, these CNN-based approaches contribute little to improving message recovery accuracy.
To address this issue, other works [Kishore et al.(2021)Kishore, Chen, Wang, Li, and Weinberger, Chen et al.(2023)Chen, Kishore, and Weinberger] employ offline optimization, directly optimizing the stego image itself by iteratively adding adversarial perturbations to it. We refer to this approach as the iterative method, in contrast with the previous approaches, termed non-iterative methods. While this iterative approach can improve message hiding accuracy, it requires additional optimization steps for each image during both training and inference, leading to an inference time exceeding 600 ms per image and significantly compromising the imperceptibility.
Inspired by the difficulties encountered in previous works, we now turn our attention to the details of the message encoding procedure in these CNN-based models. We reveal their main constraint: CNNs disorder the message order, thus having limited accuracy. Thus, differing from previous CNN-based methods, we are the first to leverage MLPs, MHSA and PEs to design an order-preserving framework.
Multi-head Self-Attention. Attention mechanism is widely adopted in the domain of Natural Language Processing [Vaswani et al.(2017a)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez, Kaiser, and Polosukhin] and various computer vision tasks [Carion et al.(2020)Carion, Massa, Synnaeve, Usunier, Kirillov, and Zagoruyko, Zhu et al.(2020)Zhu, Su, Lu, Li, Wang, and Dai, Ranftl et al.(2021)Ranftl, Bochkovskiy, and Koltun, Bai et al.(2022)Bai, Hu, Zhu, Huang, Chen, Fu, and Tai, Li et al.(2022)Li, Yu, Meng, Caine, Ngiam, Peng, Shen, Lu, Zhou, Le, et al.]. The benefits of it stem from the dynamic weights and long-range encoding capability [Khan et al.(2022)Khan, Naseer, Hayat, Zamir, Khan, and Shah]. We adopt the W-MSA [Liu et al.(2022)Liu, Hu, Lin, Yao, Xie, Wei, Ning, Cao, Zhang, Dong, et al.] as the fundamental building block of the message-image fusion, enabling long contextual fusion of message and image features with less computational complexity.
Figure 2: The overview of StegaFormer  consists of two main pipelines: (a) Message Concealment and (b) Message Recovery, which are facilitated by three proposed components: (c) Global Message-Image Fusion (GMIF), (d) Order-Preserving Message Encoder (OPME), and (e) Order-Preserving Message Decoder (OPMD). The sizes indicated above the modules represent the shapes of their respective outputs. LN denotes Layer Normalization.
@@ 3 Methods
[…]
@@ 3.1.1 Message Layout Preparation
A secret message $M\in\left\{0,1\right\}^{L}$, consisting of $L$ bits, is structured in the form $M\in\left\{0,1\right\}^{H\times W\times D}$ in previous approaches [Zhang et al.(2019a)Zhang, Cuesta-Infante, and Veeramachaneni, Tan et al.(2021)Tan, Liao, Liu, Cao, and Jiang, Yin et al.(2023)Yin, Wu, Wang, Lu, Zhou, and Huang, Chen et al.(2023)Chen, Kishore, and Weinberger]. Here, $H$ and $W$ represent the spatial dimensions of the message tensor, which correspond to the height and width of the cover image, while $D$ denotes the message capacity (e.g., $D=1$ for 1 BPP message capacity). Although this message layout can be easily processed using convolutional layers, it does introduce inappropriate limitations, as discussed in Sec. 1.
To address the issue, we introduce an unique message layout to introducing sequential order in the raw message bits. In this message layout, the message $M\in\left\{0,N_{r}\right\}^{L}$ is represented as $M_{\text{seg}}\in\left\{0,N_{r}\right\}^{N_{\text{ms}}\times L_{\text{ms}}}$ where $N_{\text{ms}}\times L_{\text{ms}}=L$. Here, $N_{r}$ denotes the range of message elements and $N_{\text{ms}}$ represents the number of message segments, $L_{\text{ms}}$ is the length of each message segments, and $L_{\text{ms}}\ll L$. Unless stated otherwise, $N_{r}$ is set to 1, as in previous approaches. We further explore expanding $N_{r}$ from 0 to 1 to higher ranges, such as 0 to 3 (representing 2 bits per message element), for high message capacity modes. Considering the multi-level structure of StegaFormer, we configure $N_{\text{ms}}$ and $L_{\text{ms}}$ at different sizes to align with the shape of the corresponding image features.
@@ 3.1.2 Image Feature Preparation
Our model adopt 2D convolution as image feature extractor. Given a cover image $I_{\text{cover}}\in\mathbb{R}^{H\times W\times 3}$, where $H$ and $W$ represent the height and width of the cover image, we employ two stages of operations to extract image feature $F_{\text{im}}\in\mathbb{R}^{(H/2^{i}\times W/2^{i})\times 2^{i}\cdot C}$. Here, $i=1,2$ indicates the number of stages, and $C$ is the number of channels in the output feature. For models with different message capacities, we set $C=2\times L_{\text{ms}}$. Each stage consists of 2D convolution layers followed by a down-sampling operation. The down-sampling operation applies convolution to reduce the spatial size of the feature map by half while doubling its feature depth. The output of the image feature preparation phase is $F_{\text{im}}\in\mathbb{R}^{(H/4\times W/4)\times 4C}$.

### v5

# Effective Message Hiding with Order-Preserving Mechanisms

@@ Abstract
Message hiding, a technique that conceals secret message bits within a cover image, aims to achieve an optimal balance among message capacity, recovery accuracy, and imperceptibility. While convolutional neural networks have notably improved message capacity and imperceptibility, achieving high recovery accuracy remains challenging. This challenge arises because convolutional operations struggle to preserve the sequential order of message bits and effectively address the discrepancy between these two modalities. To address this, we propose StegaFormer, an innovative MLP-based framework designed to preserve bit order and enable global fusion between modalities. Specifically, StegaFormer incorporates three crucial components: Order-Preserving Message Encoder (OPME), Decoder (OPMD) and Global Message-Image Fusion (GMIF). OPME and OPMD aim to preserve the order of message bits by segmenting the entire sequence into equal-length segments and incorporating sequential information during encoding and decoding. Meanwhile, GMIF employs a cross-modality fusion mechanism to effectively fuse the features from the two uncorrelated modalities. Experimental results on the COCO and DIV2K datasets demonstrate that StegaFormer surpasses existing state-of-the-art methods in terms of recovery accuracy, message capacity, and imperceptibility. Our code is released in https://github.com/boschresearch/Stegaformer.
@@ 1 Introduction
Image steganography is a technique that conceals confidential information within a publicly accessible image  Mielikainen (2006); Pevnỳ et al. (2010); Zhang et al. (2019a); Tan et al. (2021). In this study, we focus on message hiding. It has a wide range of applications such as copy-right protection, light field messaging Wengrowski and Dana (2019), virtual reality and augmented reality applications Tancik et al. (2020) and face anonymization Kishore et al. (2021). The goal of it is embedding secret message bits within a cover image to produce a new image (called stego image). Ensuring a balanced trade-off among message capacity, recovery accuracy, and imperceptibility is the task’s core. Generally, message capacity is measured by bits per pixel (BPP). Recovery accuracy indicates the message fidelity after being processed by these methods. Imperceptibility is minimizing the perceptible differences between the cover image and the stego image.
@@
Figure 1: The illustration of our proposed methods. (a) traditional CNN-based message encoding and (b) our proposed MLP-based message encoding: Order-Preserving Message Encoder (OPME) compares the difference in message encoding. (c) traditional feature fusion and (d) our proposed Global Message-Image Fusion (GMIF) is the difference in message-image fusion.
@@
Traditional message hiding methods often rely on heuristic approaches, such as the Least Significant Bit method Mielikainen (2006), which modifies individual pixels of a digital image to conceal message bits. Recently, with the advancements of deep learning, latest methods have significantly improved the message capacity and imperceptibility Zhang et al. (2019a); Yu (2020); Tan et al. (2021); Wei et al. (2022); Yin et al. (2023). These methods involve aligning and concatenating message bits with the cover image in a spatial manner while utilizing Convolutional Neural Network (CNN) to seamlessly integrate the message and image.
@@
However, CNN-based approaches often yield suboptimal results in message encoding and recovery, particularly when employing convolutional operations to encode multiple layers of message bits as illustrated in Fig.˜1(a). We have identified two main limitations in current CNN-based methods. Firstly, the indiscriminate use of convolutional kernels to encode message bits disrupts the sequential order of the original message bits. Secondly, the element-wise concatenation of message and image features, followed by convolutional operations, underestimates the substantial disparity between these two modalities, resulting in inadequate fusion as shown in Fig.˜1(c).
To validate these assumptions, we gradually modified the kernel size of SteganoGAN Zhang et al. (2019a)  and designed experiments at a 4 BPP message capacity using Div2k Agustsson and Timofte (2017) dataset. As depicted in Fig.˜2, reducing the kernel size enhances the model’s message recovery accuracy. With a decrease in kernel size from 7 to 1, 2D convolution is simplified to linear layer, focusing on 1D message segments along the channel dimension of the message tensor and encoding these segments sequentially. This process maintains the sequential order of the original bit string. On the other hand, the stego image quality improves as the kernel size increases due to a larger receptive field introduces more global interactions between the message and image features.
Figure 2: The change in message recovery accuracy and stego image quality was investigated across different kernel sizes using the traditional CNN-based method SteganoGAN. The results reveal that the CNN-based structure struggles to enhance message recovery accuracy while maintaining imperceptibility.
To address these challenges, as shown in Fig.˜1(b), we leverage Multi-Layer Perceptrons (MLP) to design a more appropriate method for message encoding. The MLP-based message encoding phase can focus on small and complete message segments individually. Later, the Multi-Head Self-Attention (MSA) Vaswani et al. (2017b) and positional embeddings (PEs) help to globally encode the message features while explicitly adding sequential information to the encoded message features. For message and image fusion, as shown in Fig.˜1(d), we use Windowed Multi-Head Self-Attention Layer (W-MSA) to address the limitation of receptive fields in the CNN-based structure, enabling global interactions between message and image features.
Our approach, named StegaFormer, consists of three proposed modules for enhanced message encoding, decoding, and message-image fusion, respectively. We design the first two modules, Order-Preserving Message Encoder (OPME) and Order-Preserving Message Decoder (OPMD), to effectively handle message segments while maintaining their order. The OPMD follows the same methodology and has a symmetric structure with OPME to decode the message. Long message bits are segmented into equal-length short messages. Later, MLPs encode and decode them separately O’shea and Hoydis (2017). Additionally, OPME and OPMD take full advantage of sequential order by incorporating PEs into the message features and globally encodes and decodes the features using MSA. The last one, Global Message-Image Fusion (GMIF) employs a W-MSA to fuse the information from the secret message and cover image to the stego image globally, greatly minimizing perceptible artifacts in the stego image.
Comprehensive experiments conducted on real-world datasets demonstrate the superior performance of our model in terms of recovery accuracy, message capacity, and imperceptibility. Importantly, our method exhibits a significant improvement in recovery accuracy. For message capacities ranging from 1 to 4 BPP, StegaFormer maintains a message recovery accuracy exceeding $99\%$. Even for high-capacity messages, StegaFormer achieves a message recovery accuracy of over $96\%$ at 6 BPP, surpassing the state-of-the-art method that achieves a 3 BPP capacity but with an even lower recovery accuracy. Our contributions are summarized as follows:
  • We are the first to point out that CNN disorders message features and is not appropriate for message hiding; consequently, we transfer to an MLP-based approach design.
  • We introduce two novel modules OPME & OPMD, which enhance the message encoding and decoding process by incorporating sequential order into the message.
  • We design GMIF, which leverages the global interaction of W-MSA to effectively fuse secret message and cover image features to stego image features.
  • Experimental results on real-world datasets demonstrate the remarkable superiority of our framework over existing state-of-the-art methods in terms of accuracy, capacity, and imperceptibility.
Figure 3: The overview of StegaFormer  consists of two main pipelines: (a) Message Concealment and (b) Message Recovery, which are facilitated by three proposed components: (c) Global Message-Image Fusion (GMIF), (d) Order-Preserving Message Encoder (OPME), and (e) Order-Preserving Message Decoder (OPMD). The sizes indicated above the modules represent the shapes of their respective outputs. LN denotes Layer Normalization, and W-MSA denotes Windowed Multi-Head Self-Attention Layer.
2 Related Work
2.1 Message Hiding
Message hiding aims to embed confidential message bits within a cover image, resulting in a stego image, and recover message bits from the stego image. Of course, it also needs to minimize any perceptible differences between cover image and stego image. Thus, the research history is around how to balance the three properties: message capacity, recovery accuracy, and imperceptibility.
Traditional methods often involve altering individual pixels in an image using heuristic techniques. They generally have a high recovery accuracy and imperceptibility, but low capacity. Specifically, early studies Mielikainen (2006); Pevnỳ et al. (2010) manipulate the least significant bits of the cover image. Others Chandramouli et al. (2003); Bi et al. (2007); Reddy and Raja (2009); Holub and Fridrich (2012) extend this approach to the frequency domain by transforming the cover image to different types of frequency domains for better performance. Despite their capacity for precise secret message recovery, these methods can only hide an extremely low BPP (i.e. 0.4) to evade detection by steganalysis tools Goljan et al. (2006); Fridrich and Kodovsky (2012).
Recently, CNN-based methods for message hiding greatly improve capacity and obtain significant attention Zhang et al. (2019a); Tan et al. (2021); Chen et al. (2022); Yin et al. (2023). For instances, SteganoGAN Zhang et al. (2019a) introduces 3D message tensor aligning in the spatial dimensions and ChatGAN Tan et al. (2021) incorporates channel attention to enhance accuracy and imperceptibility meanwhile keep a high capacity. SFTN Yin et al. (2023) is proposed to address the rounding error before generating the stego image. However, these CNN-based approaches contribute little to the improvement of message recovery accuracy.
In an effort to address the issue, other works Kishore et al. (2021); Chen et al. (2023) employ offline optimization, directly optimizing the stego image itself by iteratively adding adversarial perturbations to the stego image rather than improving the message hiding model. While this approach significantly enhances message recovery accuracy, optimizing individual stego images is time-consuming, and imperceptibility is greatly compromised.
Inspired by the difficulties encountered in previous works, we now turn our attention to the details of the message encoding procedure in these CNN-based models. We reveal the main constraint of CNN-based methods: CNNs disorder the message order, thus having limited accuracy. Thus, differing from previous CNN-based methods, we are the first to leverage MLPs with PEs to design an order-preserving framework.
2.2 Multi-head Self-Attention
MSA mechanism is widely adopted in the domain of Natural Language Processing Vaswani et al. (2017a) and various computer vision tasks Carion et al. (2020); Zhu et al. (2020); Ranftl et al. (2021); Bai et al. (2022); Li et al. (2022). The benefits of it stem from the dynamic weights and long-range encoding capability Khan et al. (2022). In our approach, we adopt the W-MSA as the fundamental building block of the message-image fusion, enabling long contextual fusion of message and image features with less computational complexity.
Dataset COCO
Metric ACC PSNR SSIM
Capacity 1 BPP 2 BPP 3 BPP 1 BPP 2 BPP 3 BPP 1 BPP 2 BPP 3 BPP
@@ 2.2 Multi-head Self-Attention
[…]
Dataset DIV2K
Metric ACC PSNR SSIM
Capacity 1 BPP 2 BPP 3 BPP 1 BPP 2 BPP 3 BPP 1 BPP 2 BPP 3 BPP
[…]
Table 1: Quantitative comparisons of SteganoGAN, ChatGAN, and our method on the COCO and DIV2K datasets are presented, with the best results highlighted in bold. ACC denotes message recovery accuracy. PSNR and SSIM are two metrics to measure the level of imperceptibility.
@@ 3 Methods
As illustrated in Fig.˜3, our proposed method encompasses two pipeline for message concealment, i.e., Fig.˜3(a), and message recovery, i.e., Fig.˜3(b), respectively. This approach sets itself apart from previous works by incorporating three crucial components: the GMIF, as depicted in Fig.˜3(c), for the fusion between features from the secret message and cover image, the OPME, as shown in Fig.˜3(d), for order-preserving message encoding, OPMD as shown in Fig.˜3(e) for corresponding message decoding.
@@ 3.1 Preparation
In the message concealment, cover image and secret message are pre-processed before the message-image interaction.
@@ 3.1.1 Message Layout Preparation
A secret message $M\in\left\{0,1\right\}^{L}$, consisting of $L$ bits, is structured in the form $M\in\left\{0,1\right\}^{H\times W\times D}$ in previous approaches Zhang et al. (2019a); Tan et al. (2021); Chen et al. (2022); Yin et al. (2023). Here, $H$ and $W$ represent the spatial dimensions of the message tensor, which correspond to the height and width of the cover image, while $D$ denotes the message capacity (e.g., $D=1$ for 1 BPP message capacity). Although this message layout can be easily processed using convolutional layers, it does introduce inappropriate limitations, as discussed in Sec.˜1.
To address the issue, we introduce an unique message layout, where the message $M\in\left\{0,N_{r}\right\}^{L}$ is reshaped into $M_{\text{seg}}\in\left\{0,N_{r}\right\}^{N_{\text{ms}}\times L_{\text{ms}}}$. Here, $N_{r}$ denotes the range of message elements (unless explicitly stated otherwise, $N_{r}$ is set to $1$) and $N_{\text{ms}}$ represents the number of message segments, $L_{\text{ms}}$ represents the length of each message segments, and $L_{\text{ms}}\ll L$. Considering the multi-level structure of StegaFormer, we configure $N_{\text{ms}}$ and $L_{\text{ms}}$ at different sizes to align with the shape of the corresponding image features.
@@ 3.1.2 Image Feature Preparation
Given a cover image $I_{\text{cover}}\in\mathbb{R}^{H\times W\times 3}$, where $H$ and $W$ represent the height and width of the cover image, we employ two stages of operations to extract image feature $F_{\text{im}}\in\mathbb{R}^{(H/2^{i}\times W/2^{i})\times 2^{i}\cdot C}$. Here, $i=1,2$ indicates the number of stages, and $C$ is the number of channels in the output feature. For models with different message capacities, we set $C=2\times L_{\text{ms}}$. Each stage consists of 2D convolution layers followed by a down-sampling operation. The down-sampling operation applies convolution to reduce the spatial size of the feature map by half while doubling its feature depth. The output of the image feature preparation phase is $F_{\text{im}}\in\mathbb{R}^{(H/4\times W/4)\times 4C}$.